In the formula L=3/4 bxh, b=2, x=7, and h= 1/2, L equals what?
Answer by Rollerprincess(4)

1. plug the numbers into the equation. L = 3/4 (2 * 7 * 1/2)
2. solve the part in parentheses (2 * 7 * 1/2) = 14 * 12 = 7
3. L = 3/4 of 7. 3/4 of 7 = 21/4 = 5.25.
4. L = 5.25